# Analysis of the $q\bar{q}\to Z^{*}\to hA\to 4\tau$ process within the lepton-specific 2HDM at the LHC

### v2

@@ IV Collider phenomenology
In this section, we present a detailed MC analysis for both signal and background events at the detector level. In the $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ process, i.e., the light Higgs (pseudo)scalar decays into two $\tau$ leptons and one of the $\tau$’s decays into a charged lepton and neutrino, while the other $\tau$ decays hadronically (thus labeled as $\tau_{h}$).
The main SM backgrounds are: $t\overline{t}\to l\nu_{l}l\nu_{l}b\overline{b}$, $W^{\pm}tb\to l\nu_{l}l\nu_{l}b\overline{b}$ (where the $W^{\pm}$ boson and the $b$ jet are not from a top quark resonance), $WWjj\to l\nu_{l}l\nu_{l}jj$, $Zjj\to l\nu_{l}l\nu_{l}jj$, $ZZ\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$, $t\overline{t}Z\to l\nu_{l}l\nu_{l}b\overline{b}\tau\tau$, $t\overline{t}ZZ\to l\nu_{l}l\nu_{l}l\nu_{l}b\overline{b}\tau\tau$ and $t\overline{t}WW\to l\nu_{l}l\nu_{l}l\nu_{l}jj$. To suppress the huge SM background process $Zjj$, we deliberately choose two SS leptons plus two hadronic $\tau$’s (for signal events, this is about $10\%$ of the total number of $4\tau$ ones). As a result, the final state consists of two SS leptons and two (light) jets.
@@ IV.1 Event generation and selection
We use MadGraph5-aMC@NLO v3.4.0 Alwall:2014hca to calculate cross sections and generate both signal and background events at the parton level. The following acceptance cuts are adopted for the signal and background generations:
@@
The parton-level events are then passed to Pythia Sjostrand:2006za ; Sjostrand:2014zea to simulate initial- and final-state radiation (i.e. the QED and QCD emissions), parton shower, hadronisation and heavy flavour decays. We further use Delphes v3.4.2 deFavereau:2013fsa to simulate the detector effects. For each event, the anti-$k_{t}$ jet algorithm Cacciari:2008gp is used to cluster the jets with the jet parameter $\Delta R=0.4$ in the FastJet package Cacciari:2011ma . The following kinematic cuts are further adopted in order to emulate the detector acceptance:
@@
Then, we select a final state consisting of two leptons and two jets from detector-simulated events. Tau tagging is not applied to improve the signal event rate. The signal results are presented in Tab. 3 and the relevant background processes are summarised in Tab. 4. The number of background events appears to be approximately $10^{5}$ larger than the number of signal events, with the dominant background processes at this stage $pp\to t{\bar{t}}$ and $pp\to Zjj$.
@@
The cross sections of the signal and background processes after the detector acceptance cuts, the selection condition assuming a $2l2j$ final state, and plus the two SS leptons requirement are shown in Tabs. 3 and  4, respectively.
@@
In Fig. 3, the transverse momentum distributions of the (second-)leading jets (frame (a)) and leptons (frame (b)) are presented for the signal (e.g. BP4) and all background events. It is clear that the transverse momenta of the signal final state are always much softer than those for background events: This is because they emerge from rather light Higgs bosons, whose energies are capped at masses of the order of 100 GeV.
@@ IV.2 Event reconstruction
To further improve the signal-to-background discrimination, we reconstruct some kinematic features of signal events. However, it is extremely difficult to completely reconstruct the momentum of each light Higgs state. This is because each Higgs boson decays into two $\tau$, with one of the $\tau$’s decaying to a lepton and two neutrinos plus two $\tau$ neutrinos emerging from hadronic $\tau$ decays, leading to the presence of six neutrinos in a signal event.
@@
$\chi^{2}={(M_{{lj}}^{1}-M_{{h}})}^{2}+{(M_{lj}^{2}-M_{A})}^{2},$ (8)
where $M_{h(A)}$ is the mass of the $h$($A$) (pseudo)scalar. We pair the two leptons and two jets to find a combination that minimises it and then assign these to $M_{lj}^{1}$ and $M_{lj}^{2}$, respectively. Since we did not consider neutrinos during such a mass reconstruction, the final $M_{lj}^{1,2}$ values should be smaller than the real $M_{h/A}$ ones; see Figs. 4 frames (a) and (b). Taking again BP4 as an example, the reconstructed light Higgs mass distributions are shown in Fig. 4. A weak positive correlation was observed between these two variables (with $r$ value of 0.27), as exemplified in Fig. 4 frame (c).
Figure 4: Distributions in invariant mass of a lepton and a jet for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. The reconstructed values, corresponding to the best $\chi^{2}$ fit, are labeled as $M_{lj}^{1}$ and $M_{lj}^{2}$, which are normalised in (a) and (b), respectively, and unnormalised (for $L~{}=~{}\text{300}~{}\text{fb}^{-1}$) when correlated in (c). In the bottom frame the signal (background) is in blue(red).
In order to further separate the signal from background events, we can also combine the two leptons and two jets, denoting their invariant masses as $M_{ll}$ and $M_{jj}$, respectively. For some SM background processes, these lepton and/or jet pairs always come from some heavy resonance, i.e.. a top (anti)quark and a $W^{\pm}$ or $Z$ boson, the background spectra should have peaks in special mass regions. However, such a feature is largely removed when selecting SS leptons, as shown in Fig. 5, so it cannot really be exploited here.
@@
Figure 7: Normalised distributions in missing transverse energy (a) and $M_{\text{T2}}$ (b) for BP4 and backgrounds at the LHC with $\sqrt{s}~{}=~{}\text{14 TeV}$ and $L~{}=~{}\text{300}~{}\text{fb}^{-1}$. In the right frame we show the three classes of background identified in the text separately.
We can further reconstruct the visible invariant mass, i.e. the one made up of the two leptons and two jets, $M_{\rm vis}$, but without any MET. In the signal, these particles come from a virtual resonance $Z^{*}$, hence $M_{\rm vis}$ should be smaller than in the backgrounds, where this is not normally the case, the latter thus yelding a larger value; see Fig. 6 frame (a). Another useful kinematic variable is $H_{T}$, which is the scalar sum of the transverse momenta of the final-state objects, shown in Fig. 6 frame (b). These two observables are generated in a very similar way between signal and background, which is evident by comparing the two frames in this figure.
In our process, two neutrinos come from leptonic $\tau$ decays, hence very soft because the $\tau$ leptons are from cascade decays. There are also neutrinos in the backgrounds from $W^{\pm}$, $Z$ boson or top (anti)quark (semi)leptonic decays, which have momenta higher than those of the neutrinos in signal events. The MET distribution is shown in Fig. 7 (a). Furthermore, the $M_{\text{T2}}$ variable is often used to track the masses of unseen pairs of particles Lester:1999ma ; Barr:2003ma , which is defined as:
@@
We thus reconstruct the $M_{\text{T2}}$ variable for both signal and background, while splitting the latter into three classes: top (including $t\bar{t}$, $t\bar{t}W^{+}W^{-}$, $t\bar{t}Z$, $t\bar{t}ZZ$ events), $W^{\pm}$ boson (including $W^{+}W^{-}jj$, $W^{\pm}tb$ events) and $Z$ boson (including $ZZ$ events). The $M_{\text{T2}}$ distributions are shown in Fig. 7 (b). It can be seen that those of the signal and $Z$ boson background are below 100 GeV or so while they are close to 150 GeV or so for the top quark and $W^{\pm}$ boson processes. Thus, the $M_{\text{T2}}$ variable provides an efficient discrimination, at least, against the top quark and $W^{\pm}$ boson noises.
@@ IV.3 Multivariate analysis
To optimise the signal and background distinction, we exploit a ML analysis, that is, a Gradient-Boosted Decision Tree (GBDT) approach is further applied, which is implemented in the Toolkit for Multivariate Date Analysis (TMVA) with ROOT Therhaag:2010zz .
A total of ten input variables are used in such a GBDT/TMVA analysis, as listed in Tab. 5. In addition to the invariant masses mentioned above, $M_{\text{T2}}$, and $H_{T}$, we also compute five angular variables between pairs of final-state objects: $\cos(\theta_{l_{1}j_{1}})$, $\cos(\theta_{l_{1}j_{2}})$, $\cos(\theta_{l_{2}j_{1}})$, $\cos(\theta_{l_{2}j_{2}})$. These angles are defined as the cosine of the angle between the two final states leptons $l$ and jets $j$, ordered in $E_{T}$, and $\cos(\theta_{lj-lj})$ is the angle between two reconstructed bound states $(lj)^{1}$ and $(lj)^{2}$. The distributions are shown in Fig. 8. As seen here, these variables are expected to be particularly useful since, in the signal process, they tend to be collinear due to the underlying boost structure, unlike in the background case.
Energy variables $M_{lj}^{1}$ $M_{lj}^{2}$ $M_{ll}$ $M_{jj}$ $H_{T}$ $M_{\text{T2}}$
@@
In the training stage, we take the signal and $ZZ$ background events as input. Because their physical process are very similar and difficult to distinguish apart in the selection with kinematic cuts. The $M_{\text{T2}}$ and the invariant mass $M_{lj}^{1,2}$ are the most important variables while training. But the angular variables $\cos(\theta_{lj-lj})$ and $\cos(\theta_{l_{i}j_{k}}),~{}i\neq k$ are also useful. Ultimately, for the GBDT output, a very good separation between the signal and background can be achieved, which is shown in Fig. 9.
Figure 9: GBDT output for BP4 and backgrounds.
@@ IV.4 Significances
The process of searching for signals at high-energy colliders using ML techniques is ultimately a statistical process. Therefore, to look for a signal, a large number of events need to be collected, and the kinematic variables of the final-state particles need to be analysed to find differences from the noise, but then a sizeable statistical sample needs to be collected, so as to have enough significance, in order to claim sensitivity to the signal which is being searched for.
The latter is defined here as $\Sigma=\frac{N_{S}}{\sqrt{(N_{S}+N_{B})}}$, where $N_{S}$ and $N_{B}$ represent the number of events for the signal ($B$) and (total) background ($B$), respectively, at a given luminosity $L$.
To distinguish the signal from the backgrounds and to compute the ensuing significance, we apply six kinematic cuts on reconstructed observables and one cut on the GBDT output, the details of which are listed in Tab. 6 (for BP4 in the signal case) and where $L~{}=~{}\text{300}~{}\text{fb}^{-1}$ (with a centre-of-mass energy of 14 TeV, as usual). Following the acceptance cuts and upon applying the SS lepton criterion (the combination of which we refer to as preselection, hereafter), we see a strong decrease of the initially huge SM backgrounds to the same order as the signal. For example, the troublesome $Zjj$ is almost eliminated after this cut (thus, it is no longer included in this table). As Figs. 46 show, the kinematic selection is already quite efficient in separating the signal from the background, especially for the variable $M_{\text{T2}}$. Finally, we also applied a GBDT discrimination before computing the final significances.
@@
$HT\in[60,400]~{}\text{GeV}$ 492.8 2743.7 260.6 682.3 3686.5 0.134 7.62
$M_{ll}\in[10,300]~{}\text{GeV}$ 490.7 2548.3 235 675.7 3459 0.142 7.81
$M_{lj}^{1}\in[0,120]~{}\text{GeV}$ 434.8 1656.5 139.9 486.5 2282.9 0.190 8.34
$M_{lj}^{2}\in[0,150]~{}\text{GeV}$ 415.6 1490.1 123.1 457.6 2070.8 0.20 8.34
$M_{vis}\in[80,500]~{}\text{GeV}$ 364.7 950.5 88.7 393.1 1432.3 0.255 8.6
$MT^{2}\in[10,110]~{}\text{GeV}$ 293.8 0 0 149.5 149.5 1.965 13.96
${\rm GBDT}\in[-0.35,1]$ 247.8 0 0 39.6 56.8 4.363 14.2
@@
The results on $\Sigma$ for all 9 BPs are shown in Tab. 7. With only the acceptance and SS lepton criteria, the significances can reach 3. After applying the kinematic cuts, including exploiting the $M_{\text{T2}}$ observable, the significances can be even larger than 10. Upon finally applying the GBDT selection, $\Sigma$ values can increase even further (albeit only slightly). In short, it is clear that a judicious combination of preselection and kinematic cuts combined with a GBDT analysis can result in large sensitivity to a significant expanse of 2HDM Type-X parameter space through our chosen signal at the current LHC stage, already.
@@
After kinematic cuts, w/o $M_{\text{T2}}$ GBDT 8.31 7.46 10.71 10.14 8.15 7.91 7.93 10.87 7.78
With $M_{\text{T2}}$, w/o GBDT 11.76 10.81 14.38 13.96 11.54 11.33 11.37 14.65 11.24
With $M_{\text{T2}}$ GBDT 11.91 10.85 14.23 14.20 11.94 11.57 11.52 14.31 11.12
@@
V Conclusions

### v3

@@ IV Collider phenomenology
In this section, we present a detailed MC analysis for both signal and background events at the detector level. In the $pp\to Z^{*}\to hA\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$ process, i.e., the light Higgs (pseudo)scalar decays into two $\tau$ leptons and one of the $\tau$’s decays into a charged lepton and neutrino, while the other $\tau$ decays hadronically (thus labelled as $\tau_{h}$).
In the CMS experiment, there are two search strategies for light scalars in term of topologies of final states and different topologies have different SM backgrounds. The first strategy targets a light Higgs boson h (or a pseudoscalar A) with a mass in the range of 5-15 GeV, produced in pairs from the decay process $H\to hh$ CMS:2015twz ; CMS:2017dmg ; CMS:2019spf . Subsequently, each light Higgs boson decays into two tau leptons, with at least one tau lepton decaying into a muon. Since the light scalars from Higgs decay are highly boosted, the 4 tau final states might form a topology with a pair of two collimated tau decay products. Thus, the backgrounds should include Drell–Yan di-muon pairs associated with jets, WW+jets, WW, WZ, ZZ, and QCD multi-jets. In the second strategy, the light scalar bosons are assumed to be in the mass range of 15–62.5 GeV CMS:2017dmg ; CMS:2018ams and the final states with $2\mu 2\tau$, $2\mu 2b$ and $4\mu$ are analyzed. The four objects in the final states are well separated. For these final states of signals, the irreducible $ZZ$ background, and reducible ZZ+jets, $t\bar{t}$ and WZWZ+jets backgrounds CMS:2017dmg should be taken into account. In this work, the light Higgs boson and the pseudoscalar are around 100 GeV and they are not highly boosted from their production processes and the majority of their decay products are well-separated. Therefore, we primarily adopt the analysis of the second strategy. Thus, we will consider the following SM backgrounds: $t\overline{t}\to l\nu_{l}l\nu_{l}b\overline{b}$, $W^{\pm}tb\to l\nu_{l}l\nu_{l}b\overline{b}$ (where the $W^{\pm}$ boson and the $b$ jet do not originate from a top quark resonance), $WWjj\to l\nu_{l}l\nu_{l}jj$, $jj\to l\nu_{l}l\nu_{l}jj, ZZ$$Z\to l\nu_{l}l\nu_{l}\tau_{h}\tau_{h}$, $\overline{t}Z\to l\nu_{l}l\nu_{l}b\overline{b}\tau\tau, tt$$\overline{t}ZZ\to l\nu_{l}l\nu_{l}l\nu_{l}b\overline{b}\tau\tau$, and $t\overline{t}WW\to l\nu_{l}l\nu_{l}l\nu_{l}jj$.
To suppress the huge SM background process $Zjj$, we deliberately choose two same-sign (SS) leptons plus two hadronic $\tau$’s (for signal events, this is about $10\%$ of the total number of $4\tau$ ones). As a result, the final state consists of two SS leptons and two (light) jets.
In the experimental analysis of same sign lepton final states, for example CMS:2018agk , the background typically include the misidentified leptons from jets or photons, and the sign mismeasured leptons. For the misidentified leptons, due to challenging difficulties, precise information can only be estimated by using control samples of collision data. In this analysis, we can not include such typs of backgrounds. While for the sign mismeasured leptons, the rate of sign mismeasurement depends upon the region of detector, and it was found to be around $1.5\times 10^{-5}$ in the inner ECAL barrel region for instance. Thus, the contamination of some background events like $Zjj$ can be estimated, although it has a huge cross section. Considering the fact that when a realistic work point of hadronic tau tagging is 0.6 or so ATLAS:2022aip , the rejection factor of faked tau from jets can reach to 50 or above, there is no doubt that the background from sign mismeasured leptons is well under control.
@@ IV.1 Event generation and selection
We use MadGraph5-aMC@NLO v3.4.0 Alwall:2014hca to calculate cross sections and generate both signal and background events at the parton level. The following acceptance cuts are applied for signal and background generation:
@@
The parton-level events are then passed to Pythia Sjostrand:2006za ; Sjostrand:2014zea to simulate initial- and final-state radiation (i.e., QED and QCD emissions), parton showering, hadronisation and heavy flavour decays. We further use Delphes v3.4.2 deFavereau:2013fsa to simulate the detector effects. For each event, the anti-$k_{t}$ jet algorithm Cacciari:2008gp is used to cluster the jets with the jet parameter $\Delta R=0.4$ in the FastJet package Cacciari:2011ma . The following kinematic cuts are further applied in order to emulate the detector acceptance:
@@
Next, we select a final state consisting of two leptons and two hadronic tau jets from detector-simulated events. The final cut table shows that the dominant background is the irreducible $ZZ$ process, whose tau jet distribution closely resembles that of the signal. In the analysis below, in order to have a close look at the effects of kinematic observables, we leave the discussion of effects of hadronic tau tagging in the discussion section.
The signal results are presented in Tab. 3, and the relevant background processes are summarised in Tab. 4. The number of background events is approximately $10^{5}$ times larger than that of the signal events, with the dominant background processes at this stage being $pp\to t{\bar{t}}$ and $pp\to Zjj$.
@@
The cross sections of the signal and background processes after the detector acceptance cuts, the selection condition assuming a $2l2j$ final state, along with the two SS leptons requirement are shown in Tabs. 3 and 4, respectively.
@@
In Fig. 3, the transverse momentum distributions of the (second-)leading jets (frame (a)) and leptons (frame (b)) are presented for the signal (e.g., BP4) and all background events. It is clear that the transverse momenta of the signal final states are always much softer than those of background events. This is because they emerge from rather light Higgs bosons, whose energies are capped at masses of the order of 100 GeV.
@@ IV.2 Event reconstruction
To further improve the signal-to-background discrimination, we reconstruct some kinematic features of signal events. However, it is extremely difficult to completely reconstruct the momentum of each light Higgs state. This is because each Higgs boson decays into two $\tau$’s, with one of the $\tau$’s decaying to a lepton and two neutrinos, plus two $\tau$ neutrinos emerging from hadronic $\tau$ decays, leading to the presence of six neutrinos in a signal event.
@@
$\chi^{2}=\frac{{(M_{{\text{lj}}}^{1}-M_{{{h}}})}^{2}}{\sigma_{h}^{2}}+\frac{{(M_{\text{lj}}^{2}-M_{{{A}}})}^{2}}{\sigma_{A}^{2}},$ (9)
where $M_{h(A)}$ are the masses of the hh(AA) (pseudo)scalars, and $\sigma_{h(A)}$ are the uncertainties in determining the mass peak of hh(AA). In practice, it is found that when the parameters $\sigma_{h(A)}$ are taken as either the decay widths of h(A) or the experimental uncertainties, the result is almost the same. Therefore, in this analysis, we fix $\sigma_{h(A)}=1$ GeV. We pair the two leptons and two jets to find a combination that minimizes this, and then assign these to $M_{lj}^{1}$ and $M_{lj}^{2}$, respectively. Since we did not consider neutrinos during such a mass reconstruction, the final $M_{lj}^{1,2}$ values should be smaller than the real $M_{h/A}$ ones; see Figs. 4 frames (a) and (b). Taking again BP4 as an example, the reconstructed light Higgs mass distributions are shown in Fig. 4. A weak positive correlation is observed between these two variables (with a value of $r$ being 0.27), as exemplified in Fig. 4 frame (c).
Figure 4: Distributions of the invariant mass of a lepton and a jet for BP4 and background processes at the LHC with $\sqrt{s}~=~\text{14 TeV}$ and an integrated luminosity of $L=\text{300}~\text{fb}^{-1}$. The reconstructed masses, obtained from the best $\chi^{2}$ fit, are denoted by $M_{lj}^{1}$ and $M_{lj}^{2}$, and are shown in normalized form in panels (a) and (b), respectively. Panel (c) displays the unnormalized distributions (corresponding to $L=\text{300}~\text{fb}^{-1}$) with the correlation between $M_{lj}^{1}$ and $M_{lj}^{2}$. In the bottom frame, the signal (background) is in blue(red).
In order to further separate the signal from background events, we can also combine the two leptons and two jets, denoting their invariant masses as $M_{ll}$ and $M_{jj}$, respectively. For some SM background processes, these lepton and/or jet pairs always come from some heavy resonance, i.e., a top (anti)quark and a $W^{\pm}$ or $Z$ boson; the background spectra should have peaks in specific mass regions. However, such a feature is largely removed when selecting SS leptons, as shown in Fig. 5, so it cannot really be exploited here.
@@
Figure 7: Normalised distributions in missing transverse energy (a) and $M_{{T2}}$ (b) for BP4 and backgrounds at the LHC with $\sqrt{s}~=~\text{14 TeV}$ and $L~=~\text{300}~\text{fb}^{-1}$. In the right frame we show the three classes of background identified in the text separately.
We can further reconstruct the visible invariant mass, i.e. the one made up of the two leptons and two jets, $M_{\rm vis}$, but without any MET. In the signal, these particles come from a virtual resonance $Z^{*}$, hence $M_{\rm vis}$ should be smaller than in the backgrounds, where this is not normally the case; the latter thus yielding a larger value; see Fig. 6 frame (a). Another useful kinematic variable is $H_{T}$, which is the scalar sum of the transverse momenta of the final-state objects, shown in Fig. 6 frame (b). These two observables are generated in a very similar way between signal and background, which is evident by comparing the two frames in this figure.
In our process, two neutrinos come from leptonic $\tau$ decays, hence are very soft because the $\tau$ leptons are from cascade decays. There are also neutrinos in the backgrounds from $W^{\pm}$, $Z$ boson or top (anti)quark (semi)leptonic decays, which have momenta higher than those of the neutrinos in signal events. The MET distribution is shown in Fig. 7 (a). Furthermore, the $M_{{T2}}$ variable is often used to track the masses of unseen pairs of particles Lester:1999ma ; Barr:2003ma , which is defined as:
@@
We thus reconstruct the $M_{{T2}}$ variable for both signal and background, while splitting the latter into three classes: top (including $\bar{t}, tt$$\bar{t}W^{+}W^{-}$, $\bar{t}Z, tt$$\bar{t}ZZ$ events), $W^{\pm}$ boson (including $W^{+}W^{-}jj$, $W^{\pm}tb$ events) and $Z$ boson (including $ZZ$ events). The $M_{{T2}}$ distributions are shown in Fig. 7 (b). It can be seen that those of the signal and $Z$ boson background are below 100 GeV or so, while they are close to 150 GeV or so for the top quark and $W^{\pm}$ boson processes. Thus, the $M_{{T2}}$ variable provides an efficient discrimination, at least, against the top quark and $W^{\pm}$ boson backgrounds.
@@ IV.3 Multivariate analysis
To optimise the signal and background distinction, we exploit a ML analysis, that is, a Gradient-Boosted Decision Tree (GBDT) approach which is further applied, and implemented in the Toolkit for Multivariate Data Analysis (TMVA) with ROOT Therhaag:2010zz .
A total of ten input variables are used in such a GBDT/TMVA analysis, as listed in Tab. 5. In addition to the invariant masses mentioned above, $M_{{T2}}$, and $H_{T}$, we also compute five angular variables between pairs of final-state objects: $\cos(\theta_{l_{1}j_{1}})$, $\cos(\theta_{l_{1}j_{2}})$, $\cos(\theta_{l_{2}j_{1}})$, $\cos(\theta_{l_{2}j_{2}})$, and $\cos(\theta_{lj-lj})$. These angles are defined as the cosine of the angle between the two final states leptons $l$ and jets $j$, ordered in $E_{T}$, and $\cos(\theta_{lj-lj})$ is the angle between two reconstructed bound states $(lj)^{1}$ and $(lj)^{2}$. The distributions are shown in Fig. 8. As seen here, these variables are expected to be particularly useful since, in the signal process, they tend to be collinear due to the underlying boost structure, unlike in the background case.
Energy variables $M_{\text{lj}}^{1}$ $M_{\text{lj}}^{2}$ $M_{\text{ll}}$ $M_{\text{jj}}$ $H_{{T}}$ $M_{{T2}}$
@@
In the training stage, we take the signal and $ZZ$ background events as input. Because their physical process are very similar and difficult to distinguish apart in the selection with kinematic cuts. The $M_{{T2}}$ and the invariant mass $M_{\text{lj}}^{1,2}$ are the most important variables while training. But the angular variables $\cos(\theta_{lj-lj})$ and $\cos(\theta_{l_{i}j_{k}}),~i\neq k$ are also useful. Ultimately, for the GBDT output, a very good separation between the signal and background can be achieved, which is shown in Fig. 9.
In the training stage, we take the signal and $ZZ$ background events as input, because their physical processes are very similar and difficult to distinguish in the selection with kinematic cuts. The $M_{{T2}}$ and the invariant mass $M_{lj}^{1,2}$ are the most important variables during training, but the angular variables $\cos(\theta_{lj-lj})$ and $\cos(\theta_{l_{i}j_{k}}),~i\neq k$ are also useful. Ultimately, for the GBDT output, a very good separation between the signal and background can be achieved, which is shown in Fig. 9.
Figure 9: GBDT output for BP4 and backgrounds. Label: signal in black colour, but the distribution is plotted in blue
@@ IV.4 Significances
The process of searching for signals at high-energy colliders using ML techniques is ultimately a statistical process. Therefore, to look for a signal, a large number of events need to be collected, and the kinematic variables of the final-state particles need to be analysed to find differences from the noise, but then a sizeable statistical sample needs to be collected, so as to have enough significance to claim sensitivity to the signal being searched for.
The latter is defined here as $\Sigma=\frac{N_{\text{S}}}{\sqrt{(N_{\text{S}}+N_{\text{B}})}}$, where $N_{\text{S}}$ and $N_{\text{B}}$ represent the number of events for the signal (S) and (total) background (B), respectively, at a given luminosity $L$.
To distinguish the signal from the backgrounds and to compute the ensuing significance, we apply six kinematic cuts on reconstructed observables and one cut on the GBDT output, the details of which are listed in Tab. 6 (for BP4 in the signal case), where $L~=~\text{300}~\text{fb}^{-1}$ (with a centre-of-mass energy of 14 TeV, as usual). Following the acceptance cuts and upon applying the SS lepton criterion (the combination of which we refer to as preselection hereafter), we see a strong decrease in the initially huge SM backgrounds to the same order as the signal. For example, the troublesome $Zjj$ is almost eliminated after this cut (thus, it is no longer included in this table). As Figs. 46 show, the kinematic selection is already quite efficient in separating the signal from the background, especially for the variable $M_{{T2}}$. Finally, we also applied a GBDT discrimination before computing the final significances.
@@
$H_{{T}}\in[60,400]~\text{GeV}$ 492.8 2743.7 260.6 682.3 3686.5 0.134 7.62
$M_{\text{ll}}\in[10,300]~\text{GeV}$ 490.7 2548.3 235 675.7 3459 0.142 7.81
$M_{\text{lj}}^{1}\in[0,120]~\text{GeV}$ 434.8 1656.5 139.9 486.5 2282.9 0.190 8.34
$M_{\text{lj}}^{2}\in[0,150]~\text{GeV}$ 415.6 1490.1 123.1 457.6 2070.8 0.20 8.34
$M_{\text{vis}}\in[80,500]~\text{GeV}$ 364.7 950.5 88.7 393.1 1432.3 0.255 8.6
$M_{{T2}}\in[10,110]~\text{GeV}$ 293.8 0 0 149.5 149.5 1.965 13.96
${\rm GBDT}\in[-0.35,1]$ 247.8 0 0 56.8 56.8 4.363 14.2
@@
The results on $\Sigma$ for all 9 BPs are shown in Tab. 7. With only the acceptance and SS lepton criteria, the significances can reach up to 3. After applying the kinematic cuts, including exploiting the $M_{{T2}}$ observable, the significances can be even larger than 10. Upon finally applying the GBDT selection, $\Sigma$ values can increase even further (albeit only slightly). In short, it is clear that a judicious combination of preselection and kinematic cuts, combined with a GBDT analysis, can result in large sensitivity to a significant expanse of the 2HDM Type-X parameter space through our chosen signal at the current LHC stage.
@@
After kinematic cuts, w/o $M_{{T2}}$ & GBDT 6.75 5.94 9.01 8.6 6.72 6.75 6.48 9.03 6.21
With $M_{{T2}}$, w/o GBDT 11.76 10.81 14.38 13.96 11.54 11.33 11.37 14.65 11.24
With $M_{{T2}}$ & GBDT 12.2 11.0 14.62 14.20 12.26 12.0 11.80 14.66 11.66
@@
In order to have an overview on the scanned parameter space of the 2HDM, it is necessary to extend our analysis to the whole parameter space in our scan. To do such a work, it is quite demanding for computing resources which is beyond our reach. Instead, we introduce an approximate method to estimate the significance of each points in the parameter space.
In our approximate method, we use the theoretical cross sections and the detector simulation efficiencies for signal events of each point to estimate the efficiencies. We use madgraph to compute the theoretical cross sections $\sigma_{theo}$. And we utilize the results of nine benchmark points presented above to fit the detector acceptance efficiencies. Acceptance efficiencies are defined as a ratio, labeled as $\epsilon_{acc}=\sigma_{\text{sim}}/\sigma_{\text{theo}}$, where $\sigma_{sim}$ is defined as the cross section after imposing all cuts. It is found that the acceptance efficiencies is proportional to $M_{{{h}}}+M_{{{A}}}$. Physically, it reflects that heavier Higgs bosons produce harder tau leptons, which are more likely to pass the lepton and jet pt cuts. In the parameter space of our study, it is found that the typical acceptance values range from 24% to 33%, so we take the averaged value i.e. $27.5\%$, in our method. Thus we obtain the estimated signal cross section which is defined as $\sigma_{sim}=\sigma_{theo}\times\epsilon_{acc}$.
The detection efficiency describes the probability that produced particles can be correctly identified and reconstructed. By fitting the benchmark point data, it is found that the jet detection efficiency is relatively stable, which can reach to 92.3% or so. However, the lepton detection efficiency shows a slight linear dependence on the Higgs boson masses, which simply reflect the fact that heavier particles produce more energetic tau leptons. A more energetic tau lepton can sequentially decay to an energetic lepton, which is easier for detectors to detect and reconstruct, thus a better lepton detection efficiency can yield. In the mass range of this study, it is found that the lepton detection efficiency can vary from 25% to 29%.
In order to improve signal selection and to suppress backgrounds, it is necessary to undergo the kinematic analysis and event pre-selection. The efficiency of kinematic cuts is also extracted from the benchmark points presented above, and it is observed that the efficiency is negatively correlated with the total mass $M_{{{h}}}+M_{{{A}}}$ for the signal events. Using the data of all BPs, after all kinematic cuts, the dominant SM background events are from $ZZ$ process. Finally, we obtain all the fitted efficiencies, which are summarized as given below:
$\epsilon_{acc}$ $=$ $1.4764\times 10^{-3}(M_{{{h}}}+M_{{{A}}})-0.0373,$
$\epsilon_{lep}$ $=$ $8.66\times 10^{-4}\times M_{{{h}}}+0.1830,$
$\epsilon_{kin,sig}$ $=$ $-3.86\times 10^{-4}\times(M_{{{h}}}+M_{{{A}}})+0.303,$
$\epsilon_{kin,ZZ}$ $=$ $1.53\times 10^{-3}\times(M_{{{h}}}+M_{{{A}}})-1.39\times 10^{-5}M_{{{h}}}M_{{{A}}}-0.148.$ (10)
Here we would like to comment on the errors in this fitting. The dominant error arises from fitting the kinematic cut region of the signal events, and the results are dependent upon the sample events of signal and background events. Even though, with our data samples, the current results yield at most a $6\%$ discrepancy in the obtained significance, as shown in Fig. 10. In Fig.  10, we show the fluctuation of significances in our approximate method when compared with the ones from the full simulation for the nine BPs. From the results, we can conclude that our approximate method works pretty well indeed. Thus we will apply our approximate method for a rapid estimates across a large set of parameter points. The final detectable cross section can be reliably expressed as $\sigma_{detectable}=\sigma_{sim}\times\epsilon_{jet}\times\epsilon_{lep}\times\epsilon_{kin}$.
The results for the points of parameter space are shown in Fig.  11, where only maximum significances are shown in the each grid. It is observed that there exists a broad parameter space where the signal can be reachable at the LHC and the significance can reach up to 12 or higher.
Figure 10: The fluctuation of efficiencies are demonstrated for nine BPs.
Furthermore, in order to display the distribution of significance in parameter space transparently, we demonstrate our results in the form of heatmaps. In the left and right plots, we divide the parameter space into grids of $M_{{{h}}}$-$M_{{{A}}}$ and $\text{sin}(\beta-\alpha)$-$\text{tan}\beta$, respectively. For each grid cell, we take the maximum significance among allowed points as the representative value, which is shown in Fig. 11. It is obvious that these points have a high potential to be detected at the LHC.
Figure 11: Maximum significances for each grid in $M_{{{h}}}$-$M_{{{A}}}$ and $\text{sin}(\beta-\alpha)$-$\text{tan}\beta$ plane.
V Discussion and Conclusions